OpenStudy (anonymous):

what is the slope of the line x+5y=9?

OpenStudy (anonymous):

in order to find out the slope of your equation you must follow the steps: 5y = -x + 9 ; y = -1/5 * x + 9/5 . now your slope is -1/5.

OpenStudy (anonymous):

put it in the form of y=mx+b where m is the slope

OpenStudy (anonymous):

I have the equation I need the slope.

OpenStudy (anonymous):

so you must rewrite your equation in order to get the slope as MATTW20 said.

OpenStudy (anonymous):

the slope can be find when you have this form : y = mx + n . for your equation is: y = -1/5 *x + 9/5. m = -1/5 is the slope.

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

yes thanks guys!

OpenStudy (anonymous):

you re welcome. :)